OpenStudy (anonymous):

Solve.|2x + 4| - 5 = 6

OpenStudy (mertsj):

Add 5 to both sides first.

OpenStudy (anonymous):

2x+4=9

OpenStudy (anonymous):

^Correct. Adding 5 to both sides, you have |2x+4|-5+5=6+5

OpenStudy (anonymous):

|2x+4|=11 2x+4=11 Or -2x-4=11

OpenStudy (anonymous):

then divide by 2?

OpenStudy (anonymous):

2x=7, i.e. x=3.5 -2x=11+4=15 i.e. x=-7.5

OpenStudy (anonymous):

but what do i do with the 4

OpenStudy (anonymous):

Take it to the other side of the equation (as I did in both cases)

OpenStudy (anonymous):

so 2x=7

OpenStudy (anonymous):

x=3.5 okay got it! thanks!